SIR HOW CAN I FIND THE ELECTRIC FIELD AT THE TIP OF A CHARGED CONE WITH VOLUME CHARGE DENSITY RHO AND HEIGHT H AND SLENT HEIGHT L ???

Grade:12

1 Answers

Aman Bansal
13 years ago

DEAR RAVI SHARMA,

dq = 2\pi z \sigma ds = 2\pi z \sigma \sqrt{2} dz

AND THEN ON INTEGERATING the value approaches to zero...

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win

exciting gifts

by answering the questions on Discussion Forum. So help discuss any

query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar :

Click here

to download the toolbar

AMAN BANSAL